OpenStudy (anonymous):

Find the second term in this sequence: tn= (6-n)^2

OpenStudy (ybarrap):

$$ \Large t_n=(6-n)^2 $$ 1st term is \(t_0\), second term is \(t_1\): $$ t_1=(6-1)^2=5^2=25 $$ Make sense?

OpenStudy (anonymous):

yes thankyou!!

OpenStudy (perl):

its not clear if n starts at n=1 or n = 0

OpenStudy (ybarrap):

Typically, indices start at 0, unless it's forbidden, as in 1/n.

OpenStudy (loser66):

@ybarrap no, it doesn't show to me. \[\text{No, it doesn't show}\]

OpenStudy (ybarrap):

If you can't see the LaTex, follow these instructions @Loser66 - http://openstudy.com/study#/updates/52de57e0e4b003c643a10b24